OpenStudy (anonymous):

Hi I am super new to this, I thought I would give it a shot. Can anyone help me solve this: Simplify quantity x squared plus 2 x over quantity 12 x plus 54 end quantity minus quantity 3 minus x over quantity 8 x plus 36

OpenStudy (helder_edwin):

is it like this? \[\large \frac{x^2+2x}{12x+54}-\frac{3-x}{8x+36} \]

OpenStudy (helder_edwin):

in the first fraction u can factor x in the numerator and 12 in the denominator. and in the second fraction u can factor 4 in the denominator. can u do it?

OpenStudy (helder_edwin):

sorry. i made a mistake. u cannot factor 12 from the denominator of the first fraction

OpenStudy (helder_edwin):

let's see \[\large \frac{x^2+2x}{12x+54}-\frac{3-x}{8x+36}= \frac{\color{red}{x}(x+2)}{\color{red}{6}(2x+9)}- \frac{3-x}{\color{red}{4}(2x+9)} \] the LCM is, as u said, \(12(2x+9)\)

OpenStudy (helder_edwin):

so to equal both denominators we multiply the first fraction by 2/2 and the second by 3/3

OpenStudy (helder_edwin):

\[\large \frac{2}{2}\cdot\frac{x(x+2)}{6(2x+9)}- \frac{3}{3}\cdot\frac{3-x}{4(2x+9)}= \frac{2x(x+2)}{12(2x+9)}-\frac{9-3x}{12(2x+9)}= \] \[\large =\frac{2x(x+2)\color{red}{-}(9-3x)}{12(2x+9)} \]

OpenStudy (helder_edwin):

i cannot figure out whay your teacher told u that. it makes no sense at all. maybe she got a bit confused.

OpenStudy (helder_edwin):

so to finish \[\large =\frac{2x^2+4x-9+3x}{12(2x+9)}=\frac{2x^2+7x-9}{12(2x+9)} \] factoring the numerator we got \[\large =\frac{(2x+9)(x-1)}{12(2x+9)}=\frac{x-1}{12} \]

OpenStudy (helder_edwin):

u r welcome. i log in this site whenever i have some free time. if i am not loged in, there is a lot of very intelligent people here. i am sure u will always find someone to help u.

OpenStudy (helder_edwin):

u too have a nice day. bye!